# Disruptions of stars and binary systems on chaotic orbits in the axisymmetric Milky Way center

### v1

@@ 6.5.1 Massive perturbers
We consider one last effect that could change the balance of relative rates. As argued in Perets et al. (2007), massive perturbers such as Giant Molecular Clouds (GMCs) may actually dominate the rate of scattering (which scales according to $\langle nm^{2}\rangle$) at large radii, decreasing relaxation times by potentially orders of magnitude, and increasing disruption rates by shifting regions where the loss cone is empty into the full regime.
The transition point between the empty and full loss cone regimes depends on $l_{\rm t}$, with larger values leading to a larger region where the empty loss cone regime dominates. As shown in Perets et al. (2007) and figure 12, including these massive perturbers thus has little effect on the disruption of individual stars (with small $l_{\rm t}$), but can significantly boost collisional disruption rates for giants and binary systems.
GMCs can reside in the NSD, but not in the NSC, as tidal forces would tear these relatively dilute self-gravitating bodies apart. We incorporate this into our models by applying a significant decrease to the the relaxation time around and beyond 10 pc via
@@
The equivalent version of figure 12 with the GMC-augmented relaxation rate is shown in figure 13. The rates from diving orbits are only slightly enhanced, and only for high $l_{\rm t}$ systems. At lower $l_{\rm t}$, the empty loss region does not extend to large enough radii to be affected by the change. The rates for disruption by scatterings are markedly increased (except for low $l_{\rm t}$ where systems were already in the full loss cone regime at the relevant radii). The increased scattering rate enlarges the range of radii where the full (and limiting) loss cone rate is the relevant one. Total rates are now boosted by an order of magnitude or more, and the median radii from which disrupted systems originate is now significantly larger (of order 10 pc or greater).
@@
Figure 13: The equivalent of figure 12 but now including a simple prescription for the impact on the relaxation time of massive perturbers (such as GMCs in the NSD). The previous rates (from figure 12) are shown with dashed lines (the full loss cone rates are unchanged). The diving rates have been increased by a factor of a few for larger $l_{\rm t}$ ($\gtrsim$ 10 pc2 Myr-1) while the scattering rates have increased by orders of magnitude for all but the smallest $l_{\rm t}$ shown.
As a caveat, we note that in this calculation, our use of standard relaxation time formulae implicitly assumed that the GMCs are moving isotropically. In reality, observed GC giant molecular clouds reside mostly in the disk and corotate with it. This configuration may reduce the effectiveness of GMCs at loss cone refilling (since isotropic scatterings in their frame will be tangentially biased in the GC frame). For the comparison in this section, we did not attempt to include this nuance, and only compared to the nominal effect as discussed in Perets et al. (2007). If indeed it is a flattened distribution of GMCs that dominate collisional loss cone repopulation, this may introduce an injection anisotropy into the hypervelocity star population, though of a different nature than the collisionless anisotropy we explored previously. Such collisional injection anisotropies were first found in Hamers & Perets (2017), where nuclear spiral arms were considered as the dominant massive perturber.
@@ 6.5.2 Comparing across $l_{\rm t}$
The collisionless channel is dominated by systems with lower $R_{\rm c}$ (except at $l_{\rm t}>10$). For $l_{\rm t}\lesssim 1$, i.e. values appropriate for tidal disruption of a main sequence star, most disruptions come from within $r_{\rm inf}$ and are thus expected to be marginally elliptical. In contrast, at larger $l_{\rm t}$, appropriate for the disruption of stellar giants or binary systems, most disruptions come from outside of $r_{\rm inf}$ and are expected to be marginally hyperbolic.
Whilst we cannot translate these relative rates directly into observable predictions, we can perform a very simple comparison by assuming that to first order the true rate will be $\sim f(l_{\rm t})\bar{\Gamma}$, where $f(l_{\rm t})$ is the fraction of systems in the GC with a given $l_{t}$. For example, if we were to say that the whole GC was composed of Sun-like MS stars with $l_{\rm t}\approx 0.3$ pc2 Myr-1 (such that $f(1)\sim 1$), we would predict a total TDE rate of a few $10^{-5}$ yr-1. As mentioned previously, this rate is significantly below past (collisional) TDE rate estimates made assuming isotropic velocities (Generozov et al., 2018); we find a far lower collisional rate due to tangential bias in the NSC, and this is only partially made up for by the higher collisionless rates. The tangential bias of NSC orbits on radial scales $\lesssim 10~{}{\rm pc}$ is likely also responsible for the enhancement of collisionless rates above the collisional full loss cone rate, which would not be possible for an isotropic system.
@@ 7 Discussion and Conclusions
We can call disruptions by diving orbits ’collisionless’, in contrast with the ’collisional’ disruptions of loss cone repopulation predicted from two-body scattering calculations. We compared the collisionless and collisional relative rates for our GC model, under the numerically simplifying assumption that all systems in the model have a single tidal angular momentum, $l_{\rm t}$ (i.e. the same tidal radius). We found that collisionless relative rates dominate for high $l_{\rm t}$ systems (e.g. massive evolved stars and all but the tightest binaries), often by one or more orders of magnitude. For moderate $l_{\rm t}$ (e.g. tidal disruption of a single main sequence star), the collisionless relative rates still dominate, but now only by a factor of a few, consistent with previous studies of the TDE “loss wedge” (Magorrian & Tremaine, 1999; Vasiliev & Merritt, 2013).
@@
  • The survival of high $l_{\rm t}$ systems at small radii: our calculations have focused on fixed values of $l_{\rm t}$ without acknowledging the the number of systems of a given $l_{\rm t}$ is dependent on energy (and thus radii). We might expect softer binaries (larger $l_{\rm t}$) to be collisionally ionized or driven to merger in the dense and dynamically hot inner GC, suppressing collisional and collisionless disruption rates. The peak of the rates for the collisionless mode is at smaller radii than the collisional mode, and thus if binaries of interest were strongly suppressed on scales of a few parsecs, the collisionless rate could be significantly diminished.
  • Stellar lifetimes vs the timescales of collisionless disruptions: the collisionless mode has a long associated timescale (up to 100 Myrs for systems of interest) but acts on a much larger population of systems. Comparatively the collisionless mode acts on a small population but with a shorter timescale. In balance the collisionless mode seems to dominate, but if we are interested in short-lived systems (such as B-type stars which currently are the most frequent HVS candidates Brown et al. 2018) they may not survive long enough and again the collisional rate may be diminished.
@@
  • Massive perturber populations: in this paper we have considered quite idealized models of massive perturber distributions, and in principle the collisional channel could regain dominance over collisionless binary disruptions if massive perturber populations extended to smaller radii. However, our choice of $\approx 10$ pc for the innermost GMCs is ultimately generous to the collisional channel, as this represents the tidal disruption radius for a quite extreme GMC ($\sim 10^{4}$ particles ${\rm cm}^{-3}$). The crudeness of our GMC model matters less at large radii, where the most interesting loss cone fluxes (for e.g. binaries tight enough to produce hypervelocity stars) are already “maxed out” at the full loss cone rate. Considering other massive perturbers, we note that the innermost star clusters in the MW (Arches and Quintuplet) have projected radii of $\approx 30$ pc and are thus unlikely to be important.
  • Accuracy of observed models: the collisionless rates we calculate are dominated by, and thus sensitive to, the population of NSC stars with $R_{\rm c}\sim{\rm few}\times{\rm pc}$ at $l_{z}\sim 0$. While we our chosen DF is reasonably well motivated from the most recent available data, it is still important to note that our results are most sensitive to this one feature of the underlying Galactic model.
@@ Acknowledgements
We first and foremost would like to thank Eugene Vasiliev, for their expertise and for producing and advising on the AGAMA code and its many relevant uses in this work. We also thank Mattia Sormani for discussions, references, and detailed explanations of the GC model implemented in AGAMA. We also thank Re’em Sari for multiple insightful discussions. We would also like to thank John Magorrian and Em Sandford for providing useful comments on the work. ZP and EMR acknowledge support from European Research Council (ERC) grant number: 101002511/project acronym: VEGA P. NCS gratefully acknowledges support from the Binational Science Foundation (grant Nos. 2019772 and 2020397) and the Israel Science Foundation (Individual Research Grant Nos. 2565/19 and 2414/23).
@@ A.3 NSC
It is important to note that unlike the $p_{NSD}(\rho,\lambda)$, where $b\approx 1$ seemed like a reasonable assumption, this is far from true for $p_{NSC}(\rho,\lambda)$ with the steepness decreasing across all $\rho$ (though still centred near $b=1$). This suggest that the assumption that $p_{NSC}(\epsilon,\lambda)\propto\lambda$ may not be appropriate and more care needs to be taken when using this distribution in calculations of scattering rates.

### v3

@@ 6.5.1 Massive perturbers
We consider one more effect that could change the balance of relative rates. As argued in Perets et al. (2007), massive perturbers such as Giant Molecular Clouds (GMCs) may actually dominate the rate of scattering (which scales according to $\langle nm^{2}\rangle$) at large radii, decreasing relaxation times by potentially orders of magnitude, and increasing disruption rates by shifting regions where the loss cone is empty into the full regime.
The transition point between the empty and full loss-cone regimes depends on $l_{\rm t}$, with larger values leading to a larger region where the empty loss-cone regime dominates. As shown in Perets et al. (2007) and figure 12, including these massive perturbers thus has little effect on the disruption of individual stars (with small $l_{\rm t}$), but can significantly boost collisional disruption rates for giants and binary systems.
GMCs can reside in the NSD, but not in the NSC, as tidal forces would tear these relatively dilute self-gravitating bodies apart. We incorporate this into our models by applying a significant decrease to the relaxation time around and beyond 10 pc via
@@
The equivalent version of figure 12 with the GMC-augmented relaxation rate is shown in figure 13. The rates from diving orbits are only slightly enhanced, and only for high $l_{\rm t}$ systems. At lower $l_{\rm t}$, the empty loss region does not extend to large enough radii to be affected by the change. At larger $l_{\rm t}$ the rates for disruption by scatterings are markedly increased. The increased scattering rate enlarges the range of radii where the full (and limiting) loss-cone rate is the relevant one. Total rates are now boosted by an order of magnitude or more, and the median radii from which disrupted systems originate is now significantly larger (of order 10 pc or greater).
@@
Figure 13: The equivalent of figure 12 but now including a simple prescription for the impact on the relaxation time of massive perturbers (such as GMCs in the NSD). The previous rates (from figure 12) are shown with dashed lines (the full loss-cone rates are unchanged). The diving rates have been increased by a factor of a few for larger $l_{\rm t}$ ($\gtrsim$ 10 pc2 Myr-1) while the scattering rates have increased by orders of magnitude for all but the smallest $l_{\rm t}$ shown.
As a caveat, we note that in this calculation, our use of standard relaxation time formulae implicitly assumed that the GMCs are moving isotropically. In reality, observed GC GMCs reside mostly in the disk and corotate with it. This configuration may reduce the effectiveness of GMCs at refilling depleted parameter space (since isotropic scatterings in their frame will be tangentially biased in the GC frame). For the comparison in this section, we did not attempt to include this nuance, and only compared to the nominal effect as discussed in Perets et al. (2007). If indeed it is a disk-like distribution of GMCs that dominate collisional loss-cone repopulation, this may introduce an injection anisotropy into the hypervelocity star population, though of a different nature than the collisionless anisotropy we explored previously. Such collisional injection anisotropies were first found in Hamers & Perets (2017), where nuclear spiral arms were considered as the dominant massive perturber.
@@ 6.5.2 Comparing across $l_{\rm t}$
The collisionless channel is dominated by disruptions of systems from lower $\rho$ (except at $l_{\rm t}>10$). For $l_{\rm t}\lesssim 1$, i.e. values appropriate for tidal disruption of a main sequence star, most disruptions come from within $r_{\rm inf}$ and are thus expected to be marginally elliptical. In contrast, at larger $l_{\rm t}$, appropriate for the disruption of stellar giants or binary systems, most disruptions come from outside of $r_{\rm inf}$ and are expected to be marginally hyperbolic.
[…]
Whilst we cannot translate these relative rates directly into observable predictions, we can perform a very simple comparison by assuming that to first order the true rate may be $\Gamma\sim f(l_{\rm t})\bar{\Gamma}$, where $f(l_{\rm t})$ is the fraction of systems in the GC with a given $l_{t}$. For example, if we were to say that the whole GC was composed of Sun-like MS stars with $l_{\rm t}\approx 0.3$ pc2 Myr-1 (such that $f(0.3)\sim 1$), we would predict a total TDE rate of a few $10^{-5}$ yr-1. As mentioned previously, this rate is significantly below past (collisional) TDE rate estimates made assuming isotropic velocities (Generozov et al., 2018); we find a far lower collisional rate due to tangential bias in the NSC, and this is only partially made up for by the higher collisionless rates.
@@
6.6 Comparing to Magorrian and Tremaine 1999
Our analysis is partly informed by the calculations presented in Magorrian & Tremaine (1999) (MT99), especially their Section 4, where they assess the rate of disruptions caused by precession of orbits in axisymmetric potentials. Our results broadly agree, though there are some differences to the methodology and hence the results which we thought it valuable to discuss in detail here.
Firstly, whilst MT99 also produce and draw inferences from a SoS, they only examine the state at apoapse. Unlike the SoSs used in our work, which we calculate from direct orbit integration, theirs comes from a discrete, orbit-averaged (Chirikov-like) map. This follows the method [^17] Originally this method was developed for slightly triaxial systems to model precession in the azimuthal angle, but MT99 apply it to the polar angle in the axisymmetric case. from Touma & Tremaine (1997) to calculate the successive $\theta_{\rm a},\ l_{\theta,\rm a}$ values. We have compared the two approaches (direct integration versus the orbit-averaged discrete map) and found that while the MT99 approach works well in the limit of perturbative axisymmetry, it breaks down for non-perturbative deviations from sphericity. The orbit-averaged method assumes that the variation in state between successive apoapses are small, implicitly relying on the assumption of mild axisymmetry, which is not satisfied in our GC model. They also consider regular orbits precessing into the loss wedge, though in our experiments we find that any trajectory that comes close enough to be disrupted is chaotic.
They find a critical value of $l_{\theta}$ below which orbits are assumed to be chaotic, which they term $J_{\rm l}$. To do this they perform their stepwise method on a trajectory starting from $l_{\theta,\rm a}=0$ and a range of initial $\theta_{\rm a}$. They then average over all the trajectories and take $J_{\rm l}=2\langle l_{\theta,\rm a}\rangle$. We do not find the equivalent of $J_{\rm l}$ directly, but note that it should be of order $F_{\rm dive}$ with some bias depending on the particular shape of the chaotic sea. Looking at figure 5 we can see two limitations of this approach - firstly that at apoapse not all $l_{\theta,\rm a}=0$ trajectories are chaotic, and secondly that the distribution of $l_{\theta,\rm p}$ (marginalised over $\theta$) does not in general match the distribution of $l_{\theta,\rm a}$. In general we would expect the differences between this part of their method and ours to be small, with $J_{\rm l}\sim F_{\rm dive}$ and both being used equivalently (comparing their equation 42 with our equation 57).
Our relative rates agree well with their calculation at low $l_{\rm t}$, appropriate for MS stars and hence TDEs, showing collisionless rates a factor of $\sim 2$ times larger than the collisional. The difference between the two rates is slightly larger in our results ($\sim$ a factor 10) because of the measured anisotropies of the DF, but otherwise consistent. However they do not explore larger values of $l_{\rm t}$, appropriate for the disruption of multi-body systems (and some giant stars) and it is here where the collisionless rate can dominate over the collisional one by 2 or 3 orders of magnitude (as discussed in section 6.5.2).
@@ 7 Discussion and Conclusions
We can call disruptions by diving orbits ’collisionless’, in contrast with the ’collisional’ disruptions of loss-cone repopulation predicted from two-body scattering calculations. We compared the collisionless and collisional relative rates for our GC model, under the numerically simplifying assumption that all systems in the model have a single tidal angular momentum, $l_{\rm t}$ (i.e. the same tidal radius). We found that collisionless relative rates dominate for high $l_{\rm t}$ systems (e.g. massive evolved stars and all but the tightest binaries), often by one or more orders of magnitude. For moderate $l_{\rm t}$ (e.g. tidal disruption of a single main sequence star), the collisionless relative rates still dominate, but now only by a factor of a few, consistent with previous studies of the TDE “loss wedge” (Magorrian & Tremaine, 1999; Vasiliev & Merritt, 2013).
@@
  • The survival of high $l_{\rm t}$ systems at small radii: our calculations have focused on fixed values of $l_{\rm t}$ without acknowledging the number of systems of a given $l_{\rm t}$ is dependent on energy (and thus radii). We might expect softer binaries (larger $l_{\rm t}$) to be collisionally ionized or while harder binaries may be driven to merger in the dense and dynamically hot inner GC, suppressing collisional and collisionless disruption rates. The peak of the rates for the collisionless mode is at smaller radii than the collisional mode, and thus if the binaries of interest were strongly suppressed on scales of a few parsecs, the collisionless rate could be significantly diminished.
  • Stellar lifetimes vs the timescales of collisionless disruptions: the collisionless mode has a long associated timescale (up to 100 Myrs for systems of interest) but acts on a much larger population of systems. Comparatively the collisionless mode acts on a small population but with a shorter timescale. In balance the collisionless mode seems to dominate, but if we are interested in short-lived systems (such as B-type stars which currently are the most frequent HVS candidates Brown et al. 2018) they may not survive long enough and again the collisional rate may be diminished. Similarly ongoing star formation of short lived stars may bias the DF, for example if they are mostly formed on disk-like circular orbits and need time to relax to the overall DF.
@@
  • Massive perturber populations: in this paper we have considered quite idealized models of massive perturber distributions, and in principle the collisional channel could regain dominance over collisionless binary disruptions if massive perturber populations extended to smaller radii. However, our choice of $\approx 10$ pc for the innermost GMCs is ultimately generous to the collisional channel, as this represents the tidal disruption radius for a quite extreme GMC ($\sim 10^{4}$ particles ${\rm cm}^{-3}$). The crudeness of our GMC model matters less at large radii, where the most interesting loss-cone fluxes (for e.g. binaries tight enough to produce hypervelocity stars) are already “maxed out” at the full loss-cone rate. Considering other massive perturbers, we note that the innermost star clusters in the MW (Arches and Quintuplet) have projected radii of $\approx 30$ pc and are thus unlikely to be important.
  • Accuracy of observed models: the collisionless rates we calculate are dominated by, and thus sensitive to, the population of NSC stars with $R_{\rm c}\sim{\rm few}\times{\rm pc}$ at $l_{z}\sim 0$. While we our chosen DF is reasonably well motivated from the most recent available data, it is important to note that our results are very sensitive to the anisotropy of the inner GC.
@@ Acknowledgements
We first and foremost would like to thank Eugene Vasiliev, for their expertise and for producing and advising on the AGAMA code and its many relevant uses in this work. We thank the referee for their insights and comments. We also thank Mattia Sormani for discussions, references, and detailed explanations of the GC model implemented in AGAMA. We also thank Re’em Sari for multiple insightful discussions. We would also like to thank John Magorrian and Em Sandford for providing useful comments on the work. ZP and EMR acknowledge support from European Research Council (ERC) grant number: 101002511/project acronym: VEGA P. NCS gratefully acknowledges support from the Binational Science Foundation (grant Nos. 2019772 and 2020397) and the Israel Science Foundation (Individual Research Grant Nos. 2565/19 and 2414/23).
@@ A.3 NSC
It is important to note that unlike the $p_{NSD}(\rho,\lambda)$, where $b\approx 1$ seemed like a reasonable assumption, this is far from true for $p_{NSC}(\rho,\lambda)$ with the steepness decreasing across all $\rho$ (though still centred near $b=1$). This apparent tangential bias at small radii suggests that the assumption that $p_{NSC}(\epsilon,\lambda)\propto\lambda$ may not be appropriate and more care needs to be taken when using this distribution in calculations of scattering rates.